If (x-6)^2 = 0, what is the value of 3x?
Answer by jim_thompson5910(35256)


If (x-6)^2 = 0, then x = 6 as shown by the steps below
(x-6)^2 = 0
sqrt((x-6)^2) = sqrt(0)
x-6 = 0
x-6+6 = 0+6
x = 6

We can confirm this by plugging in x = 6
(x-6)^2 = 0
(6-6)^2 = 0
(0)^2 = 0
0 = 0

Because x = 6, we can then say
3x = 3*6 = 18